Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Question

Expert replies
by chrisjim5 » Sun Oct 24, 2010 8:56 am
In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?

Can you please provide an answer to this?
Join the discussion
Source: — Problem Solving |

by shovan85 » Sun Oct 24, 2010 9:23 am
chrisjim5 wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?

Can you please provide an answer to this?
Let's the 7 people be ABCDEFG.

4 people have exactly 1 sibling:

Let A and B are siblings and C and D are siblings.

So,
A has 1 sibling B
B has 1 sibling A
C has 1 sibling D
D has 1 sibling C

3 people have exactly 2 siblings:

Let E, F and G are all siblings of each other.

So,
E has 2 siblings (F and G).
F has 2 siblings (E and G).
G has 2 siblings (E and F).

Total number of sibling pairs = 5 [AB, CD, EF, EG, FG]

Total number of pairs that can be formed from 7 people C(7,2) = 7!/5!*2! = 21

P(sibling pair) = 5/21

P(not sibling pair) = 1 - 5/21 = 16/21.
If the problem is Easy Respect it, if the problem is tough Attack it
Join the discussion

by waltz2salsa » Thu Oct 28, 2010 11:25 pm
Can someone please tell me where am i going wrong!!!

7 people,

4 have 1 siblings i.e. (a1, a2) and (b1, b2)
3 have 2 siblings i.e. (c1,c2,c3)



total ways of selecting two people out of 7 = 7c2

favorable ways:

(ways of selecting one from Ci *(ways of selecting one from left 4people) + (ways of selecting one from Ai *(ways of selecting one from left 5 people) + (ways of selecting one from Bi *(ways of selecting one from left 5 people)

= 3c1*4c1 + 2c1*5c1 + 2c1*5c1

P(no siblings) = fav ways/total ways = (3c1*4c1 + 2c1*5c1 + 2c1*5c1) / 7c2

= 2( 16/21)

:( :(
Join the discussion

by GMATGuruNY » Fri Oct 29, 2010 12:50 am
waltz2salsa wrote:Can someone please tell me where am i going wrong!!!

7 people,

4 have 1 siblings i.e. (a1, a2) and (b1, b2)
3 have 2 siblings i.e. (c1,c2,c3)



total ways of selecting two people out of 7 = 7c2

favorable ways:

(ways of selecting one from Ci *(ways of selecting one from left 4people) + (ways of selecting one from Ai *(ways of selecting one from left 5 people) + (ways of selecting one from Bi *(ways of selecting one from left 5 people)

= 3c1*4c1 + 2c1*5c1 + 2c1*5c1

P(no siblings) = fav ways/total ways = (3c1*4c1 + 2c1*5c1 + 2c1*5c1) / 7c2

= 2( 16/21)

:( :(
You're overcounting the number of non-sibling pairs:

Number of ways to combine 1 choice from (c1,c2,c3) and 1 choice from (a1,a2,b1,b2) = 3*4 = 12.

We've now counted all the non-sibling pairs that include (c1, c2, c3). We need to count only the non-sibling pairs that don't include (c1,c2,c3).
Number of ways to combine 1 choice from (a1,a2) and 1 choice from (b1,b2) = 2*2 = 4.

Total non-sibling pairs = 12+4 = 16.

Total possible pairs from 7 people = 7C2 = 21.

P(non-sibling pair) = 16/21.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion